OpenStudy (anonymous):

What is the equation of a line that passes through the point (0, 8) and is parallel to a line with a slope of 4/5?

OpenStudy (anonymous):

y = mx + b m = slope = 4/5 substitute x=0, y=8 to find b. the y-intercept

OpenStudy (anonymous):

y = 4/5x + 8?

OpenStudy (anonymous):

A.y = -4/5 x - 8 B.y = - 4/5x + 8 C.y = 4/5x - 8 D.y = 4/5x + 8

OpenStudy (anonymous):

Yes, y = (4/5)x + 8. Good.

OpenStudy (anonymous):

but u said substract x=0, y=8 so it would be -8?